Some so-called “creation scientists” claim that the Universe came into being about 6000 years ago. Find the value of the Hubble Constant H0 for such a Universe. Is this a reasonable value of H0? Explain your reasoning.
Answer #1

Relation between age of universe (T) and Hubble constant (Ho) is given as

T = 1/Ho

Therefore,

Ho = 1 / T

T = 6000 years = 6000 * 365 days * 24 hours * 60 minutes * 60 seconds = 1.8922e11 seconds

Ho = 1 / 1.8922e11

Ho = 5.285e-12

The units of Hubble constant are Km/s/Mpc

1 Mpc = 3.09e19 Km

so,

Ho = 5.285e-12 * 3.09e19

Ho = 1.63e8 Km/s/Mpc